OpenStudy (anonymous):

solve for m: 3m / 5 + a = b

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

3m/5=b-a

OpenStudy (anonymous):

3m=5b-5a

OpenStudy (anonymous):

m=(5b-5a)/3

OpenStudy (godlovesme):

Is that the question

OpenStudy (nurali):

3m / 5 + a = b 3m / 5=b-a 3m=5b-5a m=5b-5a/3

OpenStudy (anonymous):

@Godlovesme : yes

OpenStudy (godlovesme):

@nurali please donot solve the problem just walk them through thanks :) I prefer m= 5 (b-a) /3 @boricua5 still need help? :)

OpenStudy (anonymous):

is it 3/5 (b-a) ? @Godlovesme

OpenStudy (godlovesme):

No do u have choices? And how did u get that.?

OpenStudy (anonymous):

a. 5/3 (a-b) b. 5/3 (b-a) c. 3/5 (a-b) d. 3/5 (b-a)

OpenStudy (godlovesme):

Do u wAnna work out the whole thing?? Or

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whatever is easier for you .

OpenStudy (godlovesme):

Cool our original eqn was 3m/5 +a= b First subtract a It doesn't matter I want u to understand

OpenStudy (anonymous):

so i got 3m/ 5 = b-a

OpenStudy (godlovesme):

Right 3m/5 =b-a M is mmultiplied by 3 and divided by 5 do the opposite And multiply both sides by 5

OpenStudy (godlovesme):

@boricua5 what do you think the last step is?